OpenStudy (anonymous):

SOLVE PLEASE: The radius of a circular track is 210 meter. find the number of revolutions made by a man in 660 meter race?

OpenStudy (anonymous):

ALSO GIVE FORMULA

OpenStudy (anonymous):

Circumference of circle = 2*pi*radius

OpenStudy (anonymous):

every 2*pi*radius = 1 revolution

OpenStudy (anonymous):

Figured out the answer ?

OpenStudy (anonymous):

Answervis 2... but dnt know how it comes

OpenStudy (anonymous):

No, answer should be 1/2(half)

OpenStudy (anonymous):

sorry shabeer, ans = 2

OpenStudy (anonymous):

Circumference = 2*(22/7)*210 2*22*30=1320 660/1320=1/2

OpenStudy (anonymous):

Circumference = 2 * 3.14 * 210 = 1318.8m. In a 660 m race, there are 1318.8/660 rounds around the circular track.

OpenStudy (anonymous):

@ GT, we can take the value of pi to be 22/7, as there is 210, and 22/7 is more accurate than 3.14

OpenStudy (anonymous):

I agree. Exact value is less important than the arithmetic formulation involved.

OpenStudy (anonymous):

in book Ans is 2

OpenStudy (anonymous):

Then the question or answer must be wrong.

OpenStudy (anonymous):

Either the answer in the book is wrong or the problem formulation in the book is wrong.